# Ultraluminous X-ray sources in Globular Clusters

### v1

@@ 2.2 Simulations
Except dynamical evolution parameters, our simulations incorporate several key prescriptions for stars and binaries, some of them particularly relevant to ULX formation, specifically: We employed the rapid supernova mechanism as described by Fryer et al. (2012) for both neutron star and black hole formation; Neutron star and black hole natal kicks were drawn from a Maxwellian distribution with $\sigma=265\mathrel{\mathrm{km}/\mathrm{s}}$ (Hobbs et al. 2005); For black holes, these kicks were scaled by the fallback factor, which depends on the supernova prescription; Neutron stars formed through electron-capture supernovae were assigned substantially lower kicks with $\sigma=3\mathrel{\mathrm{km}/\mathrm{s}}$; Common envelope evolution was modeled using the BSE parameters $\alpha=0.5$ and $\lambda=0.0$. All simulated models in this study were set to a fixed metallicity of $\mathrm{Z}=0.001$.
More details on the MOCCA parameters can be found in Kamlah et al. (2022); Hypki et al. (2022, 2025); Giersz et al. (2024).
@@ 3.1 Number evolution of ULXs
Figure 2: Evolution of ULX numbers in globular clusters. Expected number of ULXs ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$) as a function of time since GC formation. All simulations with non-zero expected rates are presented. The line indicates the number of simulations with non-zero predictions (i.e. number of dots in this time bin). Each time bin spans $300\;\mathrel{\mathrm{Myr}}$ and the points are centered on these bins.
Figure 2 illustrates the temporal evolution of the expected number of ULXs [^3] Defined as the number of ULXs weighted by their observational probability in specific evolutionary time ranges. For example, an expected value of $0.001$ implies that, on average, one active ULX is observed across $1000$ simulations. Conversely, an expected value of $10$ indicates that, on average, $10$ ULXs are observed in a single simulation. This metric provides a probabilistic interpretation of ULX occurrence across simulations, accounting for their likelihood within evolutionary contexts. ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$) across all simulations, accompanied by the count of simulations yielding non-zero $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ for each time bin. The highest $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ is observed in the initial evolutionary stages, typically within the first $300\mathrel{\mathrm{Myr}}$ (first bin in Figure 2). This trend persists across all simulations, irrespective of parameter variations. These findings align with previous studies on field populations, where dynamical interactions are negligible, which reported that the ULX phase predominantly occurs in the early evolutionary stages of binaries (e.g., Wiktorowicz et al. 2017). Observational evidence also supports this conclusion (e.g. Wolter et al. 2018).
@@
The number of simulations with $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]\neq 0$ decreases over time, resembling predictions for populations formed in burst-like star formation episodes (e.g. Wiktorowicz et al. 2017). In GCs, this trend may result from a combination of system age effects on ULX formation (as observed in field populations) and the interplay of positive and negative effects related to dynamics. Generally, our results indicate a higher likelihood of observing ULXs in younger stellar clusters compared to older ones.
@@ 3.2 Temporal Evolution and Parameter Dependence of $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$
Figure 3: Comprehensive overview of the $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ evolution since the formation of a globular cluster (GC). Each panel represents a set of simulations with identical parameters: number of populations ($\mathrm{N}_{\mathrm{pop}}$), concentration ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$), binary fraction ($\mathrm{f}_{\mathrm{bin}}$), and the presence of new features (NF). Within each panel, ULX populations from various environments are presented: tidally filling (TF) GC, non-tidally filling (nTF) GC, nodyn (binaries evolving without dynamical interactions), and escapers (progenitors formed in a GC, but experiencing the ULX phase after ejection from the cluster). See Table 1 for simulations parameters overview. Bars and points represent the $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ in $300\mathrel{\mathrm{Myr}}$ bins
@@
The incorporation of new features (NF) has a noticeable effect on two-population simulations with moderate concentration ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}=0.2$; Figure 3, middle column). Simulations with NF exhibit higher $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ values compared to those without NF (nNF). The NF do not affect field populations as they primarily relate to the time delay for the second population.
The escaper populations generally align with the ULX population formed in situ, suggesting ongoing ejection (relacation and interactions), and thus a continuous supply of ULX progenitors to the escaper population, which later become ULXs while unbound to the cluster. There are generally fewer ULXs among escapers than in in-cluster populations. However, ULXs from escapers can dominate field populations. In general, there are no significant differences in the formation time of the ULXs from escapers to those in the field. An exception occurs in simulations with relatively strong ULX formation ($\mathrm{N}_{\mathrm{pop}}=2$, $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}=0.2$), where ULXs from escapers form at evolutionary ages not covered by field populations. For tidally filling clusters, the rate of ULX formation from escapers can exceed that of in-cluster formation, especially in simulations without new features (nNF).
@@ 3.3 Formation of ULXs
This section focuses on ULX progenitors - systems that will evolve into ULXs - and their evolution up to the initial ULX phase. It is important to note that some systems may experience multiple ULX phases interspersed with periods of quiescence or XRB phases ($\mathrm{L}_{\mathrm{X}}<10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$). Additionally, individual stars can participate in multiple ULXs, each with a different companion. This phenomenon is particularly prevalent for IMBHs (see Sect. 3.4.1).
@@
444 Number of progenitors on zams ($\mathrm{N}_{\mathrm{prog}}$), median fo the time when the ULX phase commences ($\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})$), and the total (i.e. through the entire history) expected number of ULXs ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]_{total}$). Number in parenthesis inform about the number of simulations with non-zero $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ of specific configuration. Results are present for different configurations of progenitors: A stands for the progenitor of an accretor, whereas D for the progenitor of donor, n stands for stars which do not become part of a ULX, or is a placeholder in case of single stars. p stands for pristine and means that there were no exchanges in a binary before ULX formation. First symbol represents the primary (i.e. more massive star on ZAMS), whereas the second letter the secondary. Results are divided into First and Second Populations and also between binaries and single stars. Values represent the mean and one sigma error for simulations in which the specific ULX where present.
Table 2 provides a comprehensive overview of ULX progenitors on ZAMS, i.e. stars that later became components of ULXs. The most common progenitors are those of pristine ULXs, i.e. binaries that do not undergo exchanges or disruptions prior to the ULX phase. In these systems, the primary typically becomes the accretor and the secondary the donor (A-D(p)). These progenitors evolve rapidly, initiating the ULX phase after approximately $\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})\approx 6\mathrel{\mathrm{Myr}}$, in stark contrast to other progenitors which generally begin after $\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})\gtrsim 3\mathrel{\mathrm{Gyr}}$. Pristine ULXs are observed in all simulations.
@@
Non-pristine ULXs can originate from both binary and single star progenitors on ZAMS. However, stars born in binaries more frequently become components of ULXs, even in non-pristine cases. These systems invariably undergo significant interactions, such as exchanges or disruptions, before forming a ULX. While some of these ULXs can form relatively early ($\sim 100\mathrel{\mathrm{Myr}}$), the majority emerge in later evolutionary phases ($>3\mathrel{\mathrm{Gyr}}$). In non-pristine ULXs, both the primary and secondary can act as either the accretor or donor with comparable probability. Our simulations did not reveal any non-pristine ULXs where both the primary and secondary are parts of ULXs (i.e., ”A-D” or ”D-A” type), though this is likely due to limited statistics rather than a physical constraint.
@@ 3.4 ULX properties
In this section we discuss the properties of in-cluster ULXs, i.e. these which reside inside the cluster. For discussion of ULXs among escapers see Section 3.7.
@@
555 Median properties and $95\%$ confidence intervals for ULX groups identified in this study. Column definitions: $\mathrm{k}_{\mathrm{acc}}$ accretor type; $\mathrm{k}_{\mathrm{don}}$ donor type; $\mathrm{M}_{\mathrm{acc}}$ accretor mass ($\mathrel{\mathrm{M}_{\odot}}$); $\mathrm{M}_{\mathrm{don}}$ donor mass ($\mathrel{\mathrm{M}_{\odot}}$); $a$ orbital separation ($\mathrel{\mathrm{R}_{\odot}}$); $e$ eccentricity; $\mathrm{L}_{\mathrm{X,max}}$ maximum X-ray luminosity ($\mathrel{\mathrm{erg}/\mathrm{s}}$); $\Delta\mathrm{t}_{\mathrm{ULX}}$ ULX lifetime ($\mathrel{\mathrm{Myr}}$); $\mathrm{t}_{\mathrm{phys,min}}$ ULX phase onset age. Donor types: MS - Main Sequence, HG - Hertzsprung Gap, RG - Red Giant, CHeB - Core Helium Burning. For groups with mixed donor types: * HG: $47\%$, MS: $41\%$, RG: $6\%$, CHeB: $3\%$; ** RG: $83\%$, HG: $7\%$, MS: $6\%$; *** RG: $75\%$, CHeB: $21\%$; other companion types represent less then $5\%$ of all companions.
@@
NS accretors in ULXs ($G_{\textrm{NS}}$) have median masses around $1.27\mathrel{\mathrm{M}_{\odot}}$. These systems typically have lower X-ray luminosities ($\log_{10}\mathrm{L}_{\mathrm{X,max}}\approx 39.22$) compared to their BH counterparts, but still clearly exceed the Eddington limit for a typical NS.
@@
Duty cycles vary widely among ULX groups, with $G_{\textrm{BH,MS}}$ systems showing the highest median value ($0.41$), while other groups have much lower values. This variability in duty cycles may explain the transient nature observed in many ULXs and has implications for their detectability (cf. Wiktorowicz et al. 2017).
@@
A notable subset of our sample consists of IMBH-powered systems ($G_{\textrm{IMBH}}$). These ULXs are characterized by extremely high accretor masses (median $9351\mathrel{\mathrm{M}_{\odot}}$) and the highest X-ray luminosities in our sample ($\log_{10}\mathrm{L}_{\mathrm{X,max}}\approx 40.84$). Such luminous systems, called extreme-ULXs (Wiktorowicz et al. 2019), are one of the main observational candidates for IMBHs and can provide insights into their formation and properties.

### v3

@@ 2.2 Simulations
Except dynamical evolution parameters, our simulations incorporate several key prescriptions for stars and binaries, some of them particularly relevant to ULX formation, specifically: We employed the rapid supernova mechanism as described by Fryer et al. (2012) for both neutron star (NS) and black hole (BH) formation; Neutron star and BH natal kicks were drawn from a Maxwellian distribution with $\sigma=265\mathrel{\mathrm{km}/\mathrm{s}}$ (Hobbs et al., 2005); For BHs, these kicks were scaled by the fallback factor, which depends on the supernova prescription; Neutron stars formed through electron-capture supernovae were assigned substantially lower kicks with $\sigma=3\mathrel{\mathrm{km}/\mathrm{s}}$; Common envelope evolution was modeled using the parameters from the BSE code: $\alpha=0.5$ and $\lambda=0.0$. All simulated models in this study were set to a fixed metallicity of $\mathrm{Z}=0.001$, which represents the typical metallicity of Galactic globular clusters444see the Harris catalog of Milky Way globular clusters https://physics.mcmaster.ca/~harris/mwgc.dat (Harris, 1996, 2010 update)..
@@ 3 Results
ULXs in this study are defined as binary systems with neutron star or black hole accretors exhibiting X-ray luminosities exceeding $10^{39}\text{ erg s}^{-1}$. White dwarf-powered ULXs are excluded from this analysis.
@@ 3.1 Number evolution of ULXs
Figure 2: Evolution of ULX numbers in GCs. Expected number of ULXs ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$) as a function of time since GC formation. All simulations with non-zero expected rates are presented. The line indicates the number of simulations with non-zero predictions (i.e., number of dots in this time bin). Each time bin spans $300\;\mathrel{\mathrm{Myr}}$ and the points are centered on these bins.
Figure 2 illustrates the temporal evolution of the expected number of ULXs [^5] Defined as the number of ULXs weighted by their observational probability in specific evolutionary time ranges. For example, an expected value of $0.001$ implies that, on average, one active ULX is observed across $1000$ simulations. Conversely, an expected value of $10$ indicates that, on average, $10$ ULXs are observed in a single simulation. This metric provides a probabilistic interpretation of ULX occurrence across simulations, accounting for their likelihood within evolutionary contexts. ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$) across all simulations, accompanied by the count of simulations yielding non-zero $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ for each time bin. The highest $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ is observed in the initial evolutionary stages, typically within the first $300\mathrel{\mathrm{Myr}}$ (first bin in Figure 2). This trend persists across all simulations, irrespective of parameter variations. These findings align with previous studies on field populations, where dynamical interactions are negligible, which reported that the ULX phase predominantly occurs in the early evolutionary stages of binaries (e.g., Wiktorowicz et al., 2017). Observational evidence also supports this conclusion (e.g., Wolter et al., 2018).
@@
The number of simulations with $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]\neq 0$ decreases over time, resembling predictions for populations formed in burst-like star formation episodes (e.g., Wiktorowicz et al., 2017). In GCs, this trend may result from a combination of system age effects on ULX formation (as observed in field populations) and the interplay of positive and negative effects related to dynamics. Generally, our results indicate a higher likelihood of observing ULXs in younger stellar clusters ($\lesssim 300\mathrel{\mathrm{Myr}}$) compared to older ones, which is consistent with observational results (e.g., Dage et al., 2025).
@@ 3.2 Temporal evolution and parameter dependence of $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$
Figure 3: Comprehensive overview of the $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ evolution since the formation of a GC. Each panel represents a set of simulations with identical parameters: number of populations ($\mathrm{N}_{\mathrm{pop}}$), concentration ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$), binary fraction ($\mathrm{f}_{\mathrm{bin}}$), and the presence of new features (NF). Within each panel, ULX populations from various environments are presented: tidally filling (TF) GC, non-tidally filling (nTF) GC, nodyn (binaries evolving without dynamical interactions), and escapers (progenitors formed in a GC, but experiencing the ULX phase after ejection from the cluster). See Table 7 for simulations parameters overview. Bars and points represent the $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ in $300\mathrel{\mathrm{Myr}}$ bins
@@
The incorporation of NF has a noticeable effect on two-population simulations with moderate concentration ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}=0.2$; Figure 3, middle column). Simulations with NF exhibit higher $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ values compared to those without NF. The NF do not affect field populations as they primarily relate to the time delay for the second population.
The escaper populations generally align with the ULX population formed in situ, suggesting ongoing ejection (relaxation and interactions), and thus a continuous supply of ULX progenitors to the escaper population, which later become ULXs while unbound to the cluster. There are generally fewer ULXs among escapers than in in-cluster populations. However, ULXs from escapers can dominate field populations. In general, there are no significant differences in the formation time of the ULXs from escapers to those in the field. An exception occurs in simulations with relatively strong ULX formation ($\mathrm{N}_{\mathrm{pop}}=2$, $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}=0.2$), where ULXs from escapers form at evolutionary ages not covered by field populations. For tidally filling clusters, the rate of ULX formation from escapers can exceed that of in-cluster formation, especially in simulations without NF.
@@ 3.3 Formation of ULXs
This section focuses on ULX progenitors - systems that will evolve into ULXs - and their evolution up to the initial ULX phase. It is important to note that some systems may experience multiple ULX phases interspersed with periods of quiescence or X-ray binary phases ($\mathrm{L}_{\mathrm{X}}<10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$). These phase transitions typically occur on timescales of at least hundreds of years, driven by dynamical interactions or stellar evolution processes. Additionally, individual stars can participate in multiple ULXs, each with a different companion. This phenomenon is particularly prevalent for IMBHs (see Sect. 3.5). While our models focus on these long-term changes, it’s worth noting that observations have shown both globular cluster and field ULXs can exhibit significant short-term variability on timescales of days to months (e.g., Maccarone et al., 2007; Shih et al., 2010; Brightman et al., 2023).
@@
666 Number of progenitors on ZAMS ($\mathrm{N}_{\mathrm{prog}}$), median of the time when the ULX phase commences ($\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})$), and the total (i.e., through the entire history) expected number of ULXs ($\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]_{total}$). Number in parenthesis inform about the number of simulations with non-zero $\mathbb{E}[\mathcal{N}_{\mathrm{ULX}}]$ of specific configuration. Results are present for different configurations of progenitors: A stands for the progenitor of an accretor, whereas D for the progenitor of donor, n stands for stars which do not become part of a ULX, or is a placeholder in case of single stars. p stands for pristine and means that there were no exchanges in a binary before ULX formation. First symbol represents the primary (i.e., more massive star on zero-age main sequence, or ZAMS), whereas the second letter the secondary. Results are divided into First and Second Populations and also between binaries and single stars. Values represent the mean and one sigma error for simulations in which the specific ULX where present.
Table 1 provides a comprehensive overview of ULX progenitors on zero-age main sequence (ZAMS), that is stars that later became components of ULXs. The most common progenitors are those of pristine ULXs, that is binaries that do not undergo exchanges or disruptions prior to the ULX phase, although still they can dynamically interact and change their properties. In these systems, the primary typically becomes the accretor and the secondary the donor (A-D(p)). These progenitors evolve rapidly, initiating the ULX phase after approximately $\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})\approx 6\mathrel{\mathrm{Myr}}$, in stark contrast to other progenitors which generally begin after $\mathrm{med}(\mathrm{t}_{\mathrm{ULX,start}})\gtrsim 3\mathrel{\mathrm{Gyr}}$. Pristine ULXs are observed in all simulations.
@@
Non-pristine ULXs can originate from both binary and single star progenitors on ZAMS. However, stars born in binaries more frequently become components of ULXs, even in non-pristine cases. These systems invariably undergo significant interactions, such as binary formation, exchanges or disruptions, before forming a ULX. While some of these ULXs can form relatively early ($\sim 100\mathrel{\mathrm{Myr}}$), the majority emerge in later evolutionary phases ($>3\mathrel{\mathrm{Gyr}}$). In non-pristine ULXs, both the primary and secondary can act as either the accretor or donor with comparable probability. Our simulations did not reveal any non-pristine ULXs where both the primary and secondary are parts of ULXs (i.e., ”A-D” or ”D-A” type), though this is likely due to limited statistics rather than a physical constraint.
@@ 3.4 ULX properties
In this section we discuss the properties of in-cluster ULXs, that is these which reside inside the cluster. For a discussion of ULXs among escapers see Section 3.8.
@@
777 Median properties and $95\%$ confidence intervals for ULX groups identified in this study. Column definitions: $\mathrm{k}_{\mathrm{acc}}$ accretor type (BH - black hole, NS - neutron star); $\mathrm{k}_{\mathrm{don}}$ donor type; $\mathrm{M}_{\mathrm{acc}}$ accretor mass ($\mathrel{\mathrm{M}_{\odot}}$); $\mathrm{M}_{\mathrm{don}}$ donor mass ($\mathrel{\mathrm{M}_{\odot}}$); $a$ orbital separation ($\mathrel{\mathrm{R}_{\odot}}$); $e$ eccentricity; $\mathrm{L}_{\mathrm{X,max}}$ maximum X-ray luminosity ($\mathrel{\mathrm{erg}/\mathrm{s}}$); $\Delta\mathrm{t}_{\mathrm{ULX}}$ ULX lifetime ($\mathrel{\mathrm{Myr}}$); $\mathrm{t}_{\mathrm{phys,min}}$ ULX phase onset age ($\mathrel{\mathrm{Myr}}$). Donor types: MS - main sequence, HG - Hertzsprung gap, RG - red giant, CHeB - core helium burning. For groups with mixed donor types: * HG: $47\%$, MS: $41\%$, RG: $6\%$, CHeB: $3\%$; ** RG: $83\%$, HG: $7\%$, MS: $6\%$; *** RG: $75\%$, CHeB: $21\%$; other companion types represent less then $5\%$ of all companions.
@@
Within ULXs, NS accretors ($G_{\textrm{NS}}$) have median masses around $1.27\mathrel{\mathrm{M}_{\odot}}$. These systems typically have lower X-ray luminosities ($\log_{10}\mathrm{L}_{\mathrm{X,max}}\approx 39.22$) compared to their BH counterparts, but still clearly exceed the Eddington limit for a typical NS.
@@
Duty cycles vary widely among ULX groups, with $G_{\textrm{BH,MS}}$ systems showing the highest median value ($0.41$), while other groups have much lower values. This variability in duty cycles may explain the transient nature observed in many ULXs and has implications for their detectability (cf. Wiktorowicz et al., 2017).
@@
A notable subset of our sample consists of IMBH-powered systems ($G_{\textrm{IMBH}}$). These ULXs are characterized by extremely high accretor masses (median $9351\mathrel{\mathrm{M}_{\odot}}$) and the highest X-ray luminosities in our sample ($\log_{10}\mathrm{L}_{\mathrm{X,max}}\approx 40.84$). Such luminous systems, called extreme-ULXs (Wiktorowicz et al., 2019), are one of the main observational candidates for IMBHs and can provide insights into their formation and properties.